the sum of two numbers is 44. One number is 12 more than the other. What are the numbers?
Found 2 solutions by ankor@dixie-net.com, tutorcecilia:
Answer by ankor@dixie-net.com(22740)


"One number is 12 more than the other".
one number = x
another number = (x-12)
:
"the sum of two numbers is 44."
x + (x-12) = 44
:
What are the numbers?
x + x = 44 + 12
2x = 56
x = 56/2
x = 28
:
The numbers: 28 and 16

Answer by tutorcecilia(2152)

Let the first number = x
Let the second number = x+12
The sum (addition) = 44
(x+12)+ (x) = 44
2x+12=44
2x+12-12=44-12
2x=32
2x/2=32/2
x=16
.
First number = x = 16
Second number = x+12=16+12=28
So,
16+28=44